# Data-Adaptive Graph Framelets with Generalized Vanishing Moments for Graph Signal Processing

### v2

@@ Abstract
In this paper, we propose a novel and general framework to construct tight framelet systems on graphs with localized supports based on hierarchical partitions. Our construction provides parametrized graph framelet systems with great generality based on partition trees, by which we are able to find the size of a low-dimensional subspace that best fits the low-rank structure of a family of signals. The orthogonal decomposition of subspaces provides a key ingredient for the definition of “generalized vanishing moments” for graph framelets. In a data-adaptive setting, the graph framelet systems can be learned by solving an optimization problem on Stiefel manifolds with respect to our parameterization. Moreover, such graph framelet systems can be further improved by solving a subsequent optimization problem on Stiefel manifolds, aiming at providing the utmost sparsity for a given family of graph signals. Experimental results show that our learned graph framelet systems perform superiorly in non-linear approximation and denoising tasks.
Index Terms:
@@
Graph framelets, tight frames, localized support, generalized vanishing moments, sparse representations, data-adaptive, graph signal processing, Stiefel manifolds, manifold optimization.
**footnotetext: Corresponding author
I Introduction and Motivation
@@
Graphs are prevalent data formats that have a variety of realizations in real life, e.g., social networks, traffic networks, sensor networks, etc. Given its practical importance, it is desired to have efficient signal processing tools on graphs, and the area of graph signal processing (GSP) has been one of the most active research areas during the past decade [1, 2]. Compared with signals, images, and other Euclidean-type data, graph data/signals are more irregular and flexible. Therefore, classical signal processing methods can not directly be applied to such data. In graph signal processing, a central topic is to develop notions and tools that resemble those on the continuous spaces or the discrete $d$-dimensional ($d$-D) signals, e.g., the Fourier and wavelet/framelet transforms. Stemmed from the spectral graph theory [3], the graph Fourier transforms (GFTs) based on graph Laplacians are by far one of the most fundamental concepts/tools upon which many other tools are developed.
By GFTs, notions of frequencies, low-pass filters, and high-pass filters can be defined similarly to those of the Euclidean counterparts. These concepts further facilitate the development of an analysis-synthesis framework using two-channel filtersbanks and up/down samplings on graphs analog to classical 1-D cases. The work in [4] shows that this can be established on bipartite graphs, on which up/down sampling is the simple restriction on either set of the bi-partition. This is essentially due to the algebraic property of spectral folding that is satisfied on the bipartite graphs. For arbitrary graphs, this property is not guaranteed, and one way to tackle this issue is to decompose the original graph into bipartite subgraphs. The graph signal processing is then conducted on each bipartite subgraph independently. Subsequent work improves [4] from the following aspects: (1) Using compact (polynomial) and biorthogonal filters by allowing different filter banks in the analysis and synthesis phases [5]. (2) Applying over-sampled bipartite graphs since the algorithm of decomposition in bipartite subgraphs in [4] results in a loss of edges of the original graphs [6]. (3) Improving computational complexity and quality of down samplings from using maximum spanning trees [7, 8]. (4) Generalizations to arbitrary graphs by applying generalized sampling operators, generalized graph Laplacians, and generalized inner-products [9, 10]. All of these have shown certain capabilities among various graph signal processing tasks.
Apart from the aforementioned work, there are many other realizations of graph signal analysis-synthesis tools based on various motivations and possessing other types of characteristics and merits [11, 12, 13, 14, 15, 16, 17, 18]. From a general perspective, we can unify these GSP approaches by viewing graph signals as vectors in $\mathbb{R}^{n}$ and one is aiming at the construction of representation systems for graph signals have various desired properties. The analysis and synthesis phases with respect to a representation system are typically corresponding to two matrices $\bm{T}_{a}$ and $\bm{T}_{s}$, respectively. One of the most important properties of a graph signal representation system is the perfect reconstruction property, i.e., $\bm{T}_{s}\bm{T}_{a}=\bm{I}$ with $\bm{I}$ being the identity matrix of a certain size. The property of critical sampling, i.e., $\bm{T}_{a},\bm{T}_{s}\in\mathbb{R}^{n\times n}$, is also desirable in some applications such as the graph signal compression. The perfect reconstruction allows for recovering the signal after decomposition, while the critical sampling indicates that compared with the original signal, there is no extra storage overhead after analysis.
However, critical sampling could be a drawback in some other applications. For example, in graph signal denoising, redundancy plays a crucial role in exploiting the intrinsic structures of the underlying noisy signal. By constructing systems with $\bm{T}_{a}$ as a redundant frame operator to represent graph signals rather than a non-redundant basis, one can generate smaller coefficients from such a system for noisy signals and therefore, the portions of noise have a larger chance to be thresholded. Moreover, in terms of tightness of the representation systems, except for special cases in [4, 12, 13], in the other aforementioned work, $\bm{T}_{s}\neq\bm{T}_{a}^{\top}$ (the transpose of $\bm{T}_{a}$) and thus we can not represent the analysis-synthesis framework using only $\bm{T}_{a}$. The tightness of the representation system requires the construction of $\bm{T}_{a}$ and $\bm{T}_{s}$ such that $\bm{T}_{s}=\bm{T}_{a}^{\top}$, $\bm{T}_{a}^{\top}\bm{T}_{a}=\bm{I}$ with $\bm{T}_{a}\in\mathbb{R}^{m\times n}$ for some $m\geq n$, which is equivalent to require that the rows of $\bm{T}_{a}$ forms a tight frame on $\mathbb{R}^{n}$. There are a few existing work [19, 20, 21, 22, 23] on tight frames for graph signal representations, most of which are derived from GFTs and thus are called spectral wavelets/framelets [19]. While spectral wavelets/framelets are well-interpreted in the frequency domain, the wavelets/framelets generally do not have localized support unless the filters are polynomial-type. In this case, $\bm{T}_{a}$ is a dense matrix which leads to computation and storage burden. To facilitate efficient computations, polynomial approximations to the target filters are applied. But then it fails to be a tight frame, i.e. $\bm{T}^{\top}_{a}\bm{T}_{a}\neq\bm{I}$[19]. On the other hand, following [24, 25, 26, 27] that utilize a series of hierarchical partitions on graphs, a recent work [28] proposes Haar-type graph framelets with localized support.
@@
Sparsity is one of the key characteristics for achieving efficient and effective signal processing since it implies the capability of representing signals with significantly fewer coefficients. A typical way to achieve sparsity in classical $d$-D discrete wavelet/framelet transforms is to construct high-pass filters with high vanishing moments [29, Sec. 1.2]. To elaborate, since smooth functions can be approximated by polynomials (using the Taylor expansion), polynomial sequences convolving with a filter with high vanishing moments vanish up to a certain order, and thus, the wavelet/framelet (high-pass filter) coefficients of the discrete samples of smooth functions are consequently very sparse. The theory for classical wavelet/framelet analysis is well-established; see, e.g., [30, 29, 31]. For more complicated families of signals, wavelets/framelets with more desirable properties than those of the ordinary wavelets, for example, the directionality, are required in order to achieve (optimal) sparsity for the representations of specific families of signals [32, 33, 34, 35, 36, 37, 38, 39]. On the contrary, it is generally difficult to define the corresponding notions of vanishing moments, smooth functions, and approximation properties for signals on graphs due to the irregular structure of graphs. It is also difficult to define families of signals of graphs with special properties, e.g. directional signals on graphs.
@@
Therefore, it is plausible why sparsity is seldom theoretically discussed in graph signal processing. However, there are still some attempts to define generalized vanishing moments for graphs. In [40], the property of being orthogonal to the first $k$ graph Laplacian eigenvectors is regarded as having generalized $k$ vanishing moments. A graph wavelet system is subsequently defined and the wavelets are guaranteed to have this property. For signals that are well approximated by the first $k$ eigenvectors, the system is able to produce very sparse coefficients. This inspires us to consider a more general construction of graph framelet systems in which a family of signals is not assumed to be well-approximated by the first $k$ eigenvectors but still largely lying in a $k$-dimensional subspace. Similarly, the orthogonality with this subspace can also be considered as having generalized vanishing moments and a system with this property is able to produce sparse coefficients for the family of signals. Hence, we seek to learn the low-dimensional structure of the family and construct graph framelet systems having such generalized vanishing moments. This resembles training a model in a machine learning task and can be regarded as a data-adaptive approach for the construction of graph framelet systems.
To achieve this goal, we propose a parametrized frame on graphs that generalizes previous work. As in [28], our systems are based on a partition tree that consists of a series of hierarchical partitions of the graph. Consequently, the constructed graph framelets are localized and fast transform algorithms are realizable. With the parametrization, our proposed graph framelet systems possess novel mathematical generality that includes [28] and [12] as special cases. Moreover, we propose to learn a low-dimensional subspace for a family of signals, in which we find a set of orthonormal vectors that fits a set of sample signals optimally in the $l^{2}$ sense. In such a way, our systems are expected to sparsely represent the given family of signals since the graph framelets in our system are orthogonal to this subspace. This system is further optimized so as to obtain the utmost sparsity in representing the given family of signals. The parameterization of our system is actually on the Stiefel manifolds and can be solved through two smooth optimization problems using the conjugate gradient method on the Riemannian manifolds. The graph framelets in our system thus have generalized vanishing moments and are capable of representing the given family of signals sparsely.
To summarize, the contribution of this paper is as follows:
@@
    We propose a novel, general, and fully parametrized method to construct tight frames on graphs, in which the graph framelets are with localized support and realizable fast transform algorithms.
@@
    We propose using our parameterized system to learn the low-dimensional structure of a given family of graph signals so that the resulting graph framelets possess generalized vanishing moments facilitating sparse representation.
@@
    We conduct the learning by solving two Riemannian optimization problems on the Stiefel manifolds to obtain the data-adaptive graph framelet systems.
@@
    We demonstrate the effectiveness of our proposed method by comparing it with other approaches in non-linear approximation and denoising tasks in graph signal processing.
The rest of this paper is organized as follows. The construction of $\mathcal{V}$-framelet and $\mathcal{G}$-framelet systems are described in Section II. The graph framelet transforms and their computational complexities are discussed in Section III. Data-adaptive realizations of the graph framelet systems based on Stiefel manifold optimizations are detailed in Section IV. In Section V, numerical experiments and comparisons are conducted for the tasks of non-linear approximation and denoising in graph signal processing. Conclusion and final remarks are given in Section VI.
II Graph Framelet Systems
@@
II-A Preliminaries
[…]
A countable set $\{\phi_{i}\}_{i\in\mathcal{I}}$ of a real Hilbert space $\mathcal{H}$ is called a tight frame of $\mathcal{H}$ if and only if

### v3

@@ Abstract
In this paper, we propose a general framework for constructing tight framelet systems on graphs with localized supports based on partition trees. Our construction of framelets provides a simple and efficient way to obtain the orthogonality with $k$ arbitrary orthonormal vectors. When the $k$ vectors contain most of the energy of a family of graph signals, the orthogonality of the framelets intuitively possesses “generalized ($k$-)vanishing” moments, and thus, the coefficients are sparse. Moreover, our construction provides not only framelets that are overall sparse vectors but also fast and schematically concise transforms. In a data-adaptive setting, the graph framelet systems can be learned by conducting optimizations on Stiefel manifolds to provide the utmost sparsity for a given family of graph signals. Furthermore, we further exploit the generality of our proposed graph framelet systems for heterophilous graph learning, where graphs are characterized by connecting nodes mainly from different classes. The usual assumption that connected nodes are similar and belong to the same class for homophilious graphs is contradictory for heterophilous graphs. Thus, we are motivated to bypass simple assumptions on heterophilous graphs and focus on generating rich node features induced by the graph structure, so as to improve the graph learning ability of certain neural networks in node classification. We derive a specific system of graph framelets and propose a heuristic method to select framelets as features for neural network input. Several experiments demonstrate the effectiveness and superiority of our approach for non-linear approximation, denoising, and node classification.
keywords:
@@
graph framelets, tight frames, generalized vanishing moments, data-adaptive, Stiefel manifolds, manifold optimization, node classification, heterophilous graphs
{MSCcodes}
42C40, 68T07, 65T60, 68Q06
1 Introduction
1.1 Background and Motivation
@@
Graphs are prevalent data formats that have a variety of realizations in real life, e.g., social networks, traffic networks, sensor networks, etc, and the area of graph machine learning including graph signal processing (GSP), node classification, link prediction, and so on, has been one of the most active research areas in the past decade [58, 49, 66]. Given its practical importance, it is desired to have efficient processing tools on graphs. Compared with signals, images, and other Euclidean-type data, graph data/signals are more irregular and flexible. Therefore, classical signal processing methods cannot be applied directly to such data. In graph signal processing, a central topic is to develop notions and tools that resemble those on the continuous spaces or the discrete $d$-dimensional ($d$-D) signals, e.g., the Fourier and wavelet/framelet transforms. Stemmed from the spectral graph theory [9], the graph Fourier transforms (GFTs) based on graph Laplacians are by far one of the most fundamental concepts/tools upon which many other tools are developed.
@@
Sparsity is one of the key characteristics in achieving efficient and effective signal processing since it implies the capability of representing signals with significantly fewer coefficients. A typical way to achieve sparsity in classical $d$-D discrete wavelet/framelet transforms is to construct high-pass filters with high order of vanishing moments [21, Sec. 1.2]. To elaborate, since smooth functions can be approximated by polynomials (using the Taylor expansion), polynomial sequences convolving with a filter with high order of vanishing moments vanish up to a certain order, and thus, the wavelet/framelet (high-pass filter) coefficients of the discrete samples of smooth functions are consequently sparse. The theory for classical wavelet/framelet analysis is well-established; see, e.g., [6, 11, 21, 40]. For more complicated families of signals, wavelets/framelets with more desirable properties than those of the ordinary wavelets, for example, the directionality, are required in order to achieve (optimal) sparsity for the representations of specific families of signals [3, 32, 25, 24, 77, 4, 22, 23].
@@
In contrast, it is generally difficult to define the corresponding notions of vanishing moments, smooth functions, and approximation properties for signals on graphs due to the irregular structure of graphs. It is also difficult to define families of signals of graphs with special properties, e.g., directional signals on graphs. Therefore, it is plausible that sparsity is seldom theoretically discussed in graph signal processing. However, there are still some attempts to define generalized vanishing moments for graphs. In [56], vectors with the property of being orthogonal to the first $k$ graph Laplacian eigenvectors are regarded as having generalized $k$ vanishing moments. A graph wavelet system is subsequently defined, and the wavelets are guaranteed to have this property. For signals that are well approximated by the first $k$ eigenvectors, the system is able to produce very sparse coefficients. This inspires us to consider a more general construction of graph framelet systems in which a family of signals is not assumed to be well-approximated by the first $k$ eigenvectors, but still largely lies in a $k$-dimensional subspace. Similarly, the orthogonality with this subspace can also be considered to have generalized vanishing moments, and a system with this property can produce sparse coefficients for the family of signals. More precisely, we focus on the following question:
  • Q1)
    Given a family $\mathfrak{F}$ of graph signals that is approximately included in a low-dimensional linear subspace in $\mathbb{R}^{n}$, how to define orthonormal bases and tight frames that 1) result in sparse matrices if collecting all the (frame) vectors, 2) sparsely represent vectors in $\mathfrak{F}$ by exploiting the low dimension of the linear subspace, and 3) associated with fast forward and backward transforms?
To the best of our knowledge, none of the works in the literature have considered our notion of generalized $k$-vanishing moments, and they have at least one of the following issues (see Section 1.2.1): (a) Do not allow perfect reconstruction; (b) only provide bases and thus are restricted to critical sampling; (c) result in dense matrices; and (d) without fast transforms. In this paper, we define general graph framelets that possess the property of having generalized $k$ vanishing moments and do not carry any of the above issues.
With such general graph framelet systems, we further consider other tasks in graph machine learning that could demonstrate different capabilities of our systems. We specifically consider the task of node classification on heterophilous graphs and focus on the following question:
  • Q2)
    How do we apply the general graph framelet system so that the framelets represent structural information on graphs other than the adjacency matrices?
In fact, this is motivated by the special properties of heterophilous graphs. If we inspect the classes of two nodes on a randomly chosen edge, the classes are more likely to be different. This also means that for any node, the majority of its neighbors come from other classes. This is contradictory to the usual assumption that graphs connect similar entities, upon which classical methods such as [74] utilize graph Laplacians to measure smoothness and impose smoothing on the outcomes. Thus, we are motivated to bypass simple assumptions on heterophilous graphs and focus on generating rich node features induced by the graphs. Moreover, these features should be different from the columns of the adjacency matrices, which can be considered as the most primitive structural information. One can expect that, combined with the generated features, neural networks are capable of extracting complex patterns and improving node classification on heterophilous graphs. Such an approach can be considered as feature engineering on graphs, which also appears in some of the latest GNNs (see Section 1.2.2).
Before proceeding further, we discuss some related works on graph signal processing and heterophilous graph learning below.
1.2 Related Works
1.2.1 Graph Signal Processing
By GFTs, notions of frequencies, low-pass filters, and high-pass filters can be defined similarly to those of the Euclidean counterparts. These concepts further facilitate the development of an analysis-synthesis framework using two-channel filter banks and up/down samplings on graphs analogous to classical 1-D cases. The work in [46] shows that this can be established on bipartite graphs, on which up/down sampling is the simple restriction on either set of the bipartition. This is essentially due to the algebraic property of spectral folding that is satisfied on the bipartite graphs. For arbitrary graphs, this property is not guaranteed, and one way to tackle this issue is to decompose the original graph into bipartite subgraphs. The graph signal processing is then conducted independently on each bipartite subgraph. Subsequent work improves [46] from the following aspects: (1) Using compact (polynomial) and biorthogonal filters by allowing different filter banks in the analysis and synthesis phases [47]. (2) Applying over-sampled bipartite graphs since the algorithm of decomposition in bipartite subgraphs in [46] results in a loss of edges of the original graphs [54]. (3) Improving computational complexity and quality of down-samplings by using maximum spanning trees [48, 70]. (4) Generalizations to arbitrary graphs by applying generalized sampling operators, generalized graph Laplacians, and generalized inner-products [68, 50]. All of these have shown certain capabilities among various graph signal processing tasks.
Apart from the aforementioned work, there are many other realizations of graph signal analysis-synthesis tools based on various motivations and possessing other types of characteristics and merits [38, 61, 55, 31, 44, 43, 15, 69]. From a general perspective, we can unify these GSP approaches by viewing graph signals as vectors in $\mathbb{R}^{n}$ and one is aiming at the construction of representation systems for graph signals have various desired properties. The analysis and synthesis phases with respect to a representation system are typically corresponding to two matrices $\bm{T}_{a}$ and $\bm{T}_{s}$, respectively. One of the most important properties of a graph signal representation system is the perfect reconstruction property, i.e., $\bm{T}_{s}\bm{T}_{a}=\bm{I}$ with $\bm{I}$ being the identity matrix of a certain size. The property of critical sampling, i.e., $\bm{T}_{a},\bm{T}_{s}\in\mathbb{R}^{n\times n}$, is also desirable in some applications such as the graph signal compression. The perfect reconstruction allows for recovering the signal after decomposition, while the critical sampling indicates that, compared with the original signal, there is no extra storage overhead after analysis.
However, critical sampling could be a drawback in some other applications. For example, in graph signal denoising, redundancy plays a crucial role in exploiting the intrinsic structures of the underlying noisy signal. By constructing systems with $\bm{T}_{a}$ as a redundant frame operator to represent graph signals rather than a non-redundant basis, one can generate smaller coefficients from such a system for noisy signals. Therefore, the portions of noise have a larger chance of being thresholded. Moreover, in terms of tightness of the representation systems, except for special cases in [46, 61, 55], in the other aforementioned work, $\bm{T}_{s}\neq\bm{T}_{a}^{\top}$ (the transpose of $\bm{T}_{a}$) and thus we can not represent the analysis-synthesis framework using only $\bm{T}_{a}$. The tightness of the representation system requires the construction of $\bm{T}_{a}$ and $\bm{T}_{s}$ such that $\bm{T}_{s}=\bm{T}_{a}^{\top}$, $\bm{T}_{a}^{\top}\bm{T}_{a}=\bm{I}$ with $\bm{T}_{a}\in\mathbb{R}^{m\times n}$ for some $m\geq n$, which is equivalent to require that the rows of $\bm{T}_{a}$ forms a tight frame on $\mathbb{R}^{n}$. There are a few existing works [20, 13, 2, 33, 59] on tight frames for graph signal representations, most of which are derived from GFTs and thus are called spectral wavelets/framelets [20]. While spectral wavelets/framelets are well-interpreted in the frequency domain, the wavelets/framelets generally do not have localized support unless the filters are polynomial-type. In this case, $\bm{T}_{a}$ is a dense matrix, which leads to computation and storage burdens. To facilitate efficient computations, polynomial approximations to the target filters are applied. However, it fails to be a tight frame, i.e., $\bm{T}^{\top}_{a}\bm{T}_{a}\neq\bm{I}$ [20]. On the other hand, following [10, 17, 7, 8] that utilize a series of hierarchical partitions on graphs, a recent work [36] proposes Haar-type graph framelets with localized support.
1.2.2 Node Classification in Heterophilous Graph Learning
A typical task on graph learning is to identify the classes to which each node belongs. Such classification corresponds to determining the group an account belongs to in a social network or the type of traffic connection in a traffic network. There are abundant works in the literature that deal with node classification, where graph neural network (GNN) approaches are currently one of the most active topics [66]. However, graphs are often assumed to connect similar entities. In node classification, such assumptions mean that connected nodes are more likely to be of the same classes, which is described as being homophilous. The datasets adopted in the early works are primarily homophilous, and these GNNs implicitly perform graph smoothing [30, 37]. In recent years, an ongoing active topic has been the counterpart of homophilous graphs, the so-called heterophilous graphs [71]. Intuitively, heterophily means that connected nodes are more likely to be of different classes. As a result, such a reversing nature of heterophilous graphs brings challenges to the GNNs that perform smoothing. Other types of GNNs were therefore proposed to handle node classification on heterophilous graphs.
The adaptations for heterophilous graphs in current GNNs are based on two observations: 1) Spatial aspect: For a fixed node, other nodes from the same class are distributed more outside the 1-hop neighborhood. 2) Spectral aspect: Target signals on heterophilous graphs have larger oscillations with respect to the graph Laplacian. Thus, its frequency distribution can hardly be confined to the low-frequency part. GNNs for heterophilous graphs mainly utilize graph-induced aggregations in multiple layers, which resembles the idea in deep convolutional neural networks. These GNNs can be further categorized into two types according to whether the aggregations are spatially [63, 19] or spectrally defined [12, 27, 72]. Recently, there has been a third perspective in GNNs, in which the multiple layers of neural networks are simply linear, and all graph-induced aggregations are only involved in forming inputs for the networks [42, 28]. These methods aim at jointly generating new features for all nodes. Then, with these new features as inputs, the following training is completely supervised. Such approaches have demonstrated superior performance in heterophilous node classification despite not using complicated neural network architectures. In our view, these methods can be categorized as feature engineering for heterophilous graphs.
1.3 Contributions
In the following sections, we propose framelets on graphs that generalize previous work and fit the requirements in Q1). We will also be looking at a special realization of the graph framelet system. By such an approach, we are able to provide rich, sparse graph framelets as features for node classification. Rich framelets mean that we can choose among various candidate framelets. Being sparse means that when all framelets are regarded as a single matrix, the matrix is sparse. The framelets are constructed upon an induced two-hop graph so as to provide structural information that is different from the original one-hop adjacency matrix. The framelets are then measured by the variance on the original adjacency matrix, intended for integrating one-hop and two-hop structural information. By sorting the variance, we are able to provide a spectral display of the framelets in which certain parts suggest substantial differences among the framelets. These parts are then selected as features for neural network input. The contributions of our proposed graph framelet systems are summarized as follows:
@@ 1.3 Contributions
    We propose a general method to construct tight frames on graphs, in which the graph framelets have localized support and realizable, fast, and concise transforms.
@@
    We propose using our graph framelet systems to adapt to the low-dimensional structure of a given family of graph signals so that the resulting graph framelet systems possess generalized vanishing moments, facilitating sparse representation.
@@
    We propose a special realization of the graph framelet system for generating new features as neural network input.
@@
    We demonstrate the effectiveness of our proposed method by comparing it with other approaches on non-linear approximation, denoising, and node classification.
1.4 Outline of the paper
The rest of this paper is organized as follows. The construction of $\mathcal{V}$-framelet and $\mathcal{G}$-framelet systems are described in Section 2. The graph framelet transforms and their computational complexities are discussed in Section 3. Data-adaptive realizations of the graph framelet systems based on Stiefel manifold optimizations are detailed in Section 4. In Section 5, numerical experiments and comparisons are conducted for the tasks of non-linear approximation and denoising in graph signal processing. Section 6 shows the definition of the $2$-hop graph framelet systems and their applications to node classification. Conclusion and final remarks are given in Section 7.
2 Graph Framelet Systems
@@
2.1 Preliminaries
We first introduce some necessary notation, definitions, and preliminary results. We denote an undirected weighted graph with $n$ vertices as $\mathcal{G}:=(\mathcal{V},\mathcal{E},\bm{W})$ (or simply $\mathcal{G}:=(\mathcal{V},\mathcal{E})$), where $\mathcal{V}:=\{v_{1},v_{2},\dots,v_{n}\}$, $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$, $\bm{W}=(w_{ij})_{1\leq i,j\leq n}\in\mathbb{R}^{n\times n}$ denote the set of vertices, the set of edges, and the weight (adjacency) matrix, respectively. The space $L^{2}(\mathcal{G}):=\{f\,|\,\mathcal{V}\rightarrow\mathbb{R}\}$ is the collection of graph signals on $\mathcal{G}$ and can be regarded as the finite-dimensional Hilbert space $\mathbb{R}^{n}$ with the usual Euclidean inner-product $\langle\cdot,\cdot\rangle$ and induced ($l_{2}$) norm $\|\cdot\|:=\sqrt{\langle\cdot,\cdot\rangle}$. The cardinality of a set is denoted by $|\cdot|$. Let $\bm{L}:=\bm{D}-\bm{W}$ denote the unnormalized graph Laplacian matrix, where the diagonal matrix $\bm{D}:=\text{diag}(d_{1},d_{2},\dots,d_{n})$ with $d_{i}:=\sum_{j=1}^{n}w_{ij},1\leq i\leq n$ is the degree matrix. $\bm{L}$ is positive semidefinite and thus it has $n$ eigenvalues $0\leq\lambda_{1}\leq\lambda_{2}\leq\dots\leq\lambda_{n}$ associated with real orthonormal eigenvectors $\bm{u}_{1},\bm{u}_{2},\dots,\bm{u}_{n}$. If $\mathcal{G}$ is connected, then $\bm{u}_{1}$ a vector with constant element $1/\sqrt{n}$. Let $\tilde{\bm{L}}:=\bm{I}-\bm{D}^{-1/2}\bm{W}\bm{D}^{-1/2}$ denote the normalized graph Laplacian matrix. The quadratic form $\bm{f}\tilde{\bm{L}}\bm{f}^{\top},\bm{f}\in\mathbb{R}^{1\times n}$ equals
(1) $\sum_{e=(v_{i},v_{j})\in\mathcal{E}}\frac{w_{ij}(\bm{f}_{i}-\bm{f}_{j})^{2}}{\sqrt{\bm{D}_{ii}\bm{D}_{jj}}},$
which measures the variance of a signal $\bm{f}$ on a graph by summing (normalized) value differences on all edges. For unit-norm $\bm{f}$, $\bm{f}\bm{L}\bm{f}^{\top}\in[0,2]$ [28].
For $K\in\mathbb{N}$, the set $\{1,\dots,K\}$ is denoted as $[K]$. For a matrix $\bm{M}$, we denote its $i$-th row as $\bm{M}_{i,:}$. We denote $\bm{I}$ the identity matrix of a certain size and omit its size for simplicity. In this paper, vectors in $\mathbb{R}^{d},d\in\mathbb{N}$ are assumed to be row vectors, or equivalently matrices in $\mathbb{R}^{1\times d}$. Given a set $\{\bm{X}_{i}\}_{i=1}^{m}\subseteq\mathbb{R}^{s\times t}$ of $m$ matrices, we use $\mathrm{CONCAT}(\{\bm{X}_{i}\}_{i=1}^{m}):=\llbracket\bm{X}_{i}\rrbracket_{i\in[m]}:=[\bm{X}_{1}^{\top},\ldots,\bm{X}_{m}^{\top}]^{\top}$ to denote the matrix in $\mathbb{R}^{ms\times t}$ constructed by concatenating $\bm{X}_{1},\bm{X}_{2},\dots,\bm{X}_{m}$ along the rows. We also use $\bm{\llparenthesis}\bm{X}_{i}\bm{\rrparenthesis}_{i\in[m]}:=[\bm{X}_{1},\ldots,\bm{X}_{m}]\in\mathbb{R}^{s\times mt}$ to denote concatenation along the columns. For $\bm{X}\in\mathbb{R}^{m\times d}$, we use $\mathrm{VEC}(\bm{X})$ to denote the vector in $\mathbb{R}^{md}$ constructed by concatenating the rows $\bm{X}_{i,:},1\leq i\leq m$. For $\bm{x}\in\mathbb{R}^{md}$, we use $\mathrm{MAT}(\bm{x},m)$ to denote the inverse of VEC, i.e. constructing a matrix $\bm{X}\in\mathbb{R}^{m\times d}$ from the vector $\bm{x}$ by consecutively slicing $\bm{x}$ into $m$ vectors in $\mathbb{R}^{d}$ and concatenating them along the rows. We use $\lceil\cdot\rceil$ to denote the ceiling operation of real numbers.
A countable set $\{\phi_{i}\}_{i\in\mathcal{I}}$ of a real Hilbert space $\mathcal{H}$ is called a tight frame (with frame bounds equal to $1$) of $\mathcal{H}$ if and only if